# Computational Complexity of Physical Information Sufficiency

### v3

@@ 10.7 Feature Selection
In machine learning, feature selection asks which input features are relevant for prediction. This is known to be NP-hard in general [9, 25]. Our results show the decision-theoretic analog is coNP-complete for both checking and minimization. [C:T4.6,4.7;R:S](L: CC61,CC36.)
@@ 10.9 Model Selection
Statistical model selection (AIC/BIC/cross-validation) provides practical heuristics for choosing among models [1, 56, 66]. Our results formalize the regime-level reason heuristic selection appears: without added structural assumptions, exact optimal model selection inherits worst-case intractability, so heuristic methods implement explicit weakened-guarantee policies for unresolved structure. [C:T4.6,5.1,6.1;R:E,S,S+ETH](L: CC61,CC16,CC69,CC70,CC71.)
@@ 10.11 Certifying Outputs and Proof-Carrying Claims
In the cited literature, these pillars are treated separately: representation-sensitive hardness, query-access lower bounds, and certifying soundness disciplines. This paper composes all three for the same decision-relevance object in one regime-typed and machine-checked framework. [C:T5.1;P5.3,3.16;R:S+ETH,Qf,AR](L: CC16,CC50,CC75.)
@@ 11 Conclusion
What the author contributed: The problem formulations (SUFFICIENCY-CHECK, MINIMUM-SUFFICIENT-SET, ANCHOR-SUFFICIENCY), the hardness conjectures, the tractability conditions, and the connection to over-modeling in engineering practice.
@@
This paper establishes the computational complexity of coordinate sufficiency problems:
  • SUFFICIENCY-CHECK is coNP-complete (Theorem 4.6)
  • MINIMUM-SUFFICIENT-SET is coNP-complete (Theorem 4.7)
  • ANCHOR-SUFFICIENCY is $\Sigma_{2}^{\textsf{P}}$-complete (Theorem 4.9)
  • An encoding-regime separation contrasts explicit-state polynomial-time (polynomial in $|S|$) with a succinct worst-case ETH lower bound witnessed by a hard family with $k^{*}=n$ (Theorem 5.1)
  • Full intermediate query-access lower bounds are formalized as a finite-state Opt-oracle core ($\Omega(|S|)$, Boolean instantiation $\Omega(2^{n})$) plus Boolean-interface refinements for value-entry and state-batch access, with explicit subproblem-to-full transfer, weighted randomized robustness, and oracle-lattice transfer/strictness closures (Propositions 5.35.16)
  • Tractable subcases exist for explicit-state encoding, separable utility, and tree-structured utility with explicit local factors (Theorem 6.1)
@@
[C:T4.6,4.7,4.9,5.1,6.1,2.37;P5.3,5.7,5.11,5.8,5.15,5.16,2.28,2.29,2.30,2.31,2.34;C2.35;R:E,S,S+ETH,Qf,Qb,AR] (L: CC61,CC36,CC5,CC16,CC69,CC50,L9,L8,L11,L21,L20,DQ548,DQ537,DQ539,DQ541,CT6,CT8,CT4.)
These results place the problem of identifying decision-relevant coordinates at the first and second levels of the polynomial hierarchy.
@@
The reduction constructions and key equivalence theorems are machine-checked in Lean 4 (see Appendix A for proof listings). The formalization verifies that the TAUTOLOGY reduction correctly maps tautologies to sufficient coordinate sets; complexity classifications (coNP-completeness, $\Sigma_{2}^{\textsf{P}}$-completeness) follow by composition with standard complexity-theoretic results (TAUTOLOGY is coNP-complete, $\exists\forall$-SAT is $\Sigma_{2}^{\textsf{P}}$-complete). The strengthened gadget showing that non-tautologies yield instances with all coordinates relevant is also formalized. [C:T4.6,4.7,4.9,5.1;R:S,S+ETH] (L: CC61,CC36,CC5,CC16,CC24.)
@@
Section 9 develops a quantitative consequence: when a problem requires $k$ dimensions and a model handles only $j<k$ natively, the remaining kjk-j dimensions must be handled externally at each decision site. For $n$ sites, total external work is $(k-j)\times n$.[C:T9.5;R:LC] (L: HD21.)
@@
The landscape above is complete for the static sufficiency class under C1–C4 and the declared regimes; the items below are adjacent-class extensions or secondary refinements. Several questions remain for future work:
@@
  • Physical-system instantiation: once a physical class is encoded into the declared decision interface, the same regime-typed hardness/tractability map and integrity constraints apply to that class (Section 2.7).[C:P2.34;C2.35;T2.37;R:AR](L: CT6,CT8,CT4.)
Hence the design choice is typed: enforce a tractable regime, or adopt weakened guarantees with explicit verification boundaries.[C:T6.1,5.1;R:RG](L: CC71,CC16.) Equivalently, under the attempted-competence matrix, over-specification is rational only in attempted-competence-failure cells; once exact competence is available in the active regime (or no attempted exact competence was made), persistent over-specification is irrational for the same verified-cost objective.[C:P3.60;R:AR](L: IC27.) By Proposition 3.60, the integrity-preserving matrix contains 3 irrational cells and 1 rational cell.[C:P3.60;R:AR](L: IC27.)
Epistemic Alignment Implications for Agent Training
Within this paper’s formalism, epistemic alignment is a claim-discipline property: outputs are asserted only when certifiable, and otherwise emitted as abstain reports (optionally carrying tentative guesses and self-reflected confidence under zero certified confidence). Formally, this is the integrity/competence split of Section 3, not a full value-alignment claim.[C:D3.8,3.9;P3.58;R:TR](L: IC18.)
@@
If a solver-agent declares a truth-preserving objective for claims about relation $\mathcal{R}$ in regime $\Gamma$, uncertified outputs are admissible only as $\mathsf{ABSTAIN}/\mathsf{UNKNOWN}$ rather than asserted answers. Under the signal discipline, this includes abstain-with-guess/self-confidence reports while keeping certification evidence-gated.[C:D3.19,3.20;P3.23,3.29;C3.25;R:AR](L: IC46,IC38,IC41.) Equivalently, training policies that suppress abstention in uncertified regions violate the typed claim discipline under the declared task model.[C:D3.8;P3.58,3.61,3.60;R:AR](L: IC18,CC30,IC27.)
This yields a training doctrine for solver-agents under declared tasks/regimes:
@@
    In hardness-blocked regimes, abstention or explicit guarantee-weakening is structurally required unless regime assumptions are changed.[C:T5.1;P3.61,3.60;R:S,S+ETH](L: CC16,CC30,IC27.)
Operationally, a reporting schema consistent with this framework tracks at least five quantities: integrity violations, competent non-abstaining coverage on the declared scope, abstention quality under regime shift, the calibration gap between self-reflected and certified confidence channels, and scalar runtime witness distributions (with bound-qualified fraction semantics when bounds are declared). This is consistent with bounded-rationality feasibility discipline under declared resource models [71, 5].
@@ Appendix A Lean 4 Proof Listings
The complete Lean 4 formalization is available in the companion artifact (Zenodo DOI listed on the title page). The mechanization consists of 13391 lines across 58 files, with 597 theorem/lemma statements.
@@ A.1 Lean Handle ID Map
ID Full Lean handle
@@
CC3 DecisionQuotient.ClaimClosure.agent_transfer_licensed_iff_snapshot
@@
CC6 DecisionQuotient.ClaimClosure.boundaryCharacterized_iff_exists_sufficient_subset
CC7 DecisionQuotient.ClaimClosure.bounded_actions_detectable
CC8 DecisionQuotient.ClaimClosure.bridge_boundary_represented_family
CC9 DecisionQuotient.ClaimClosure.bridge_failure_witness_non_one_step
CC10 DecisionQuotient.ClaimClosure.bridge_transfer_iff_one_step_class
CC11 DecisionQuotient.ClaimClosure.certified_total_bits_split_core
CC12 DecisionQuotient.ClaimClosure.cost_asymmetry_eth_conditional
CC13 DecisionQuotient.ClaimClosure.declaredBudgetSlice
CC14 DecisionQuotient.ClaimClosure.declaredRegimeFamily_complete
@@
CC15 DecisionQuotient.ClaimClosure.declared_physics_no_universal_exact_certifier_core
CC16 DecisionQuotient.ClaimClosure.dichotomy_conditional
CC17 DecisionQuotient.ClaimClosure.epsilon_admissible_iff_raw_lt_certified_total_core
CC18 DecisionQuotient.ClaimClosure.exact_admissible_iff_raw_lt_certified_total_core
@@
CC19 DecisionQuotient.ClaimClosure.exact_certainty_inflation_under_hardness_core
CC20 DecisionQuotient.ClaimClosure.exact_raw_eq_certified_iff_certainty_inflation_core
CC21 DecisionQuotient.ClaimClosure.exact_raw_only_of_no_exact_admissible_core
CC22 DecisionQuotient.ClaimClosure.explicit_assumptions_required_of_not_excused_core
CC23 DecisionQuotient.ClaimClosure.explicit_state_upper_core
CC24 DecisionQuotient.ClaimClosure.hard_family_all_coords_core
CC25 DecisionQuotient.ClaimClosure.horizonTwoWitness_immediate_empty_sufficient
CC26 DecisionQuotient.ClaimClosure.horizon_gt_one_bridge_can_fail_on_sufficiency
CC27 DecisionQuotient.ClaimClosure.information_barrier_opt_oracle_core
CC28 DecisionQuotient.ClaimClosure.information_barrier_state_batch_core
CC29 DecisionQuotient.ClaimClosure.information_barrier_value_entry_core
CC30 DecisionQuotient.ClaimClosure.integrity_resource_bound_for_sufficiency
CC31 DecisionQuotient.ClaimClosure.integrity_universal_applicability_core
@@
CC32 DecisionQuotient.ClaimClosure.meta_coordinate_irrelevant_of_invariance_on_declared_slice
CC33 DecisionQuotient.ClaimClosure.meta_coordinate_not_relevant_on_declared_slice
CC34 DecisionQuotient.ClaimClosure.minsuff_collapse_core
CC36 DecisionQuotient.ClaimClosure.minsuff_conp_complete_conditional
CC37 DecisionQuotient.ClaimClosure.no_auto_minimize_of_p_neq_conp
@@
CC38 DecisionQuotient.ClaimClosure.no_exact_claim_admissible_under_hardness_core
@@
CC39 DecisionQuotient.ClaimClosure.no_exact_claim_under_declared_assumptions_unless_excused_core
CC40 DecisionQuotient.ClaimClosure.no_exact_identifier_implies_not_boundary_characterized

### v4

@@ 14.8 Feature Selection
In machine learning, feature selection asks which input features are relevant for prediction. This is known to be NP-hard in general [11, 33]. Our results place the decision-theoretic analog in a regime map: static checking/minimization at coNP-complete, stochastic sufficiency at PP-complete, and sequential sufficiency at PSPACE-complete. [C:T4.6, 4.7, 6.7, 7.7;R:S, RG](L: CC61, CC36, DC9, DC10.)
@@ 14.10 Model Selection
Statistical model selection (AIC/BIC/cross-validation) provides practical heuristics for choosing among models [1, 75, 86]. Our results formalize the regime-level reason heuristic selection appears: without added structural assumptions, exact optimal model selection inherits worst-case intractability, so heuristic methods implement explicit weakened-guarantee policies for unresolved structure. [C:T4.6, 9.1, 10.1;R:E, S, S+ETH](L: CC61, CC16, CC69, CC70, CC71.)
@@ 14.12 Certifying Outputs and Proof-Carrying Claims
In the cited literature, these pillars are treated separately: representation-sensitive hardness, query-access lower bounds, and certifying soundness disciplines. This paper composes all three for the same decision-relevance object in one regime-typed and machine-checked framework. [C:T9.1;P9.3, 3.16;R:S+ETH, Qf, AR](L: CC16, CC50, CC75.)
@@ 15 Conclusion
Formally: this section composes proved regime placements, bridge theorems, and claim-typing constraints into final landscape consequences.
This paper gives a regime-first complexity map: static/stochastic/sequential placements with strict hierarchy (DC9, DC10, DC1, DC2), static inner-class decomposition by decision/access/structure layers (CCC6, CCC5, CCC1, CCC3, CCC7, CC61), theorem-level physical transport plus exact represented-adjacent bridge boundary (CT4, CC67, CC69, CC85), Bayesian uniqueness under DQ/thermodynamic admissibility axioms (FN14), physical-collapse no-go transfer chain (PH11, PH12, PH13, PH14, PH15), derived SAT-bridge map plus reduction-energy transfer and disjunctive assumption boundary (PH16, PH17, PH18, PH19, PH20, PH21, PH22, PH26, PH27), explicit finite-budget-necessity countermodels (PH31, PH32, PH33), regime-robust deterministic/probabilistic/sequential no-collapse specializations (PH28, PH29, PH30), observer-collapse anchor closure across stochastic/sequential checks (PA1, PA2, PA3, PA4, PA5, PA6, PA7, PA8, PA9), recurrent conversation-layer clamp/report bridge (CV4, CV7, CV17, CV18, CV13, CV16), explicit Neukart–Vinokur duality specialization from the thermodynamic lift (CC77, CC76), and formal centralized/distributed hardness thresholds through dominance and amortization theorems (HD1, HD4, HD3). [C:T4.6, 4.7, 4.9, 6.7, 7.7, 9.1, 10.1, 2.45, 2.54, 2.46, 3.4, 9.50, 12.19, 13.3, 13.6;P9.3, 9.28, 8.1, 8.2;C9.32;R:E, S, S+ETH, Qf, Qb, AR, RG, LC, H=cost-growth]
Informally: this conclusion combines previously proved blocks under explicit tags.
@@
What the author contributed: The informal problem formulations (SUFFICIENCY-CHECK, MINIMUM-SUFFICIENT-SET, ANCHOR-SUFFICIENCY), the informal hardness conjectures, the informal tractability conditions, and the connection to over-modeling in engineering practice.
@@
Acknowledgments
This work depends on results whose authors made possible the insights that could then be machine-verified.
Fundamental physics. Albert Einstein established special relativity and the light-speed bound [25]; Hermann Minkowski gave it the spacetime formulation [52]. Max Planck established energy quantization and the constant $k_{B}$ [68]. Rolf Landauer proved that erasure costs thermodynamic work [44]; Charles Bennett showed that reversible computation circumvents this bound [7]. Bérut et al. experimentally verified Landauer’s principle [9]. The counting gap theorem (BA10) proves that some finite $c$ must exist from pure mathematics; Einstein told us which one.
Proof infrastructure. The Lean 4 proof assistant [22] served as the arbiter of correctness. Leonardo de Moura, Sebastian Ullrich, and the Lean 4 team built a tool that transforms mathematical claims into mechanically verified theorems. The Mathlib library [88] provided the formal infrastructure: order theory, linear algebra, complexity-theoretic definitions, and the lemmas that made the formalization possible.
Methodological note. The proofs in this paper are machine-checked; their validity is independent of generation method. The human author contributed the informal problem formulations, informal hardness conjectures, informal tractability conditions, and the connection to over-modeling in engineering practice. The machine contributors (Claude, GPT-4, Lean 4) served as proof assistants, tactic explorers, and drafting partners. The final arbiter was the Lean compiler: if it accepts the proof, the theorem holds.
@@ Acknowledgments
This paper establishes a complete regime map for coordinate sufficiency and a static inner decomposition:
  • Regime map (core classes). Static, stochastic, and sequential sufficiency-check questions are placed at coNP, $P$, and PSPACErespectively, with strict static $\subset$ stochastic $\subset$ sequential separation under declared assumptions (Theorems 4.6, 6.7, 7.7; Propositions 8.1, 8.2).
  • Observer-collapse anchor closure. Under declared observer-collapse hypotheses, stochastic/sequential anchor checks are forced by effective-state singleton structure and seeded support conditions (PA1, PA4, PA5, PA6, PA7, PA9).
  • Static decision layer decomposition. Within the static class, SUFFICIENCY-CHECK and MINIMUM-SUFFICIENT-SET are coNP-complete, while ANCHOR-SUFFICIENCY is $\Sigma_{2}^{\textsf{P}}$-complete (Theorems 4.6, 4.7, 4.9).
  • Static access layer decomposition. An encoding-regime separation contrasts explicit-state polynomial-time (polynomial in $|S|$) with succinct ETH worst-case lower bounds, and query-access obstruction cores refine this with value-entry/state-batch/oracle-lattice transfers (Theorem 9.1; Propositions 9.39.16).
  • Static structure layer decomposition. Six structural subclasses provide certified polynomial recovery conditions inside the static class (Theorem 10.1).
@@
  • Exact represented-adjacent bridge boundary: transfer is licensed iff the class is one-step deterministic; represented horizon, stochastic-criterion, and transition-coupled classes have explicit bridge-failure witnesses (Theorem 2.54)
  • Exact-report admissibility is characterized by a certified-bit gap criterion; when exact reporting is blocked, exact accounting collapses to raw-only (Theorem 3.47, Corollary 3.49)
  • The thermodynamic lift yields the Neukart–Vinokur duality specialization $dU\geq\lambda\,dC$ for $\lambda>0$ (Corollary 9.32, Proposition 9.28)
[C:T4.6, 4.7, 4.9, 6.7, 7.7, 9.1, 10.1, 2.45, 2.54, 3.47;P8.1, 8.2, 9.3, 9.7, 9.11, 9.8, 9.15, 9.16, 2.34, 2.35, 2.36, 2.37, 2.42, 9.28, 3.46;C2.43, 3.49, 9.32;R:E, S, S+ETH, Qf, Qb, AR, RG, H=cost-growth] (L: CCC6, CCC5, CCC1, DC9, DC10, DC1, DC2, CCC3, CCC7, CC61, CC29, CC28, CC43, CC52, CC27, DT6, DT13, DT16, DT24, CT6, CT8, CT4, CC67, CC69, CC85, CC18, CC11, CC21, CC77, CC76.)
[…]
@@
The reduction constructions and key equivalence theorems are machine-checked in Lean 4 (see Appendix A for proof listings). The formalization verifies TAUTOLOGY and $\exists\forall$-SAT static reductions, theorem-level stochastic/sequential class placements, and strict regime-separation claims; complexity classifications (coNP, $\Sigma_{2}^{\textsf{P}}$, PP, PSPACE) follow by composition with standard complexity-theoretic results under the declared assumptions. The strengthened gadget showing that non-tautologies yield instances with all coordinates relevant is also formalized. [C:T4.6, 4.7, 4.9, 6.7, 7.7, 9.1;P8.1, 8.2;R:S, S+ETH, RG] (L: CCC6, CCC5, CCC1, DC9, DC10, CCC3, DC1, DC2.)
@@
Section 13 develops a quantitative consequence: when a problem requires $k$ dimensions and a model handles only $j<k$ natively, the remaining kjk-j dimensions must be handled externally at each decision site. For $n$ sites, total external work is $(k-j)\times n$.[C:T13.5;R:LC] (L: HD26.)
@@
The landscape above is complete for the static sufficiency class under C1–C4; the items below are adjacent-class extensions or secondary refinements. Several questions remain for future work:
@@
  • Physical-system instantiation: once a physical class is encoded into the declared decision interface, the same regime-typed hardness/tractability map and integrity constraints apply to that class (Section 2.7).[C:P2.42;C2.43;T2.45;R:AR](L: CT6, CT8, CT4.)
Physical Scope Guard.
All transfer claims in this paper are typed through explicit physical encodings, declared observables, and declared assumption maps. Transfer is admitted only on declared physical instances satisfying the physical scope gate.[C:D2.38, 2.39;P2.40, 2.42;R:AR]
Hence the design choice is typed: enforce a tractable regime, or adopt weakened guarantees with explicit verification boundaries.[C:T10.1, 9.1;R:RG](L: CC71, CC16.) Equivalently, under the attempted-competence matrix, over-specification is rational only in attempted-competence-failure cells; once exact competence is available in the active regime (or no attempted exact competence was made), persistent over-specification is irrational for the same verified-cost objective.[C:P3.60;R:AR](L: IC27.) By Proposition 3.60, the integrity-preserving matrix contains 3 irrational cells and 1 rational cell.[C:P3.60;R:AR](L: IC27.)
Measurement-Constrained Reporting for Physical Solvers
Within this paper’s formalism, reporting correctness is a claim-discipline property: outputs are asserted only when certifiable, and otherwise emitted as abstain reports (optionally carrying tentative guesses and self-reflected confidence under zero certified confidence). Formally, this is the integrity/competence split of Section 3, i.e., an evidence-typing statement for reported claims under declared regimes.[C:D3.8, 3.9;P3.58;R:TR](L: IC18.)
Formally: the abstention doctrine is a typed-report consequence of integrity/competence and evidence-gating predicates.
@@
If a physical solver system declares a truth-preserving objective for claims about relation $\mathcal{R}$ in regime $\Gamma$, uncertified outputs are admissible only as $\mathsf{ABSTAIN}/\mathsf{UNKNOWN}$ rather than asserted answers. Under the signal discipline, this includes abstain-with-guess/self-confidence reports while keeping certification evidence-gated.[C:D3.19, 3.20;P3.23, 3.29;C3.25;R:AR](L: IC46, IC38, IC41.) Equivalently, reporting protocols that suppress abstention in uncertified regions violate the typed claim discipline under the declared task model.[C:D3.8;P3.58, 3.61, 3.60;R:AR](L: IC18, CC30, IC27.)
This yields an output-control doctrine for engineered solver systems under the task/regime contract:
@@
    In hardness-blocked regimes, abstention or explicit guarantee-weakening is structurally required unless regime assumptions are changed.[C:T9.1;P3.61, 3.60;R:S, S+ETH](L: CC16, CC30, IC27.)
Operationally, a reporting schema consistent with this framework tracks at least five quantities: integrity violations, competent non-abstaining coverage on the scoped family, abstention quality under regime shift, the calibration gap between self-reflected and certified confidence channels, and scalar runtime witness distributions (with bound-qualified fraction semantics when bounds are given). This is consistent with bounded-rationality feasibility discipline under bounded resource models [92, 5].
Informally: if exact support is not certified, do not make exact claims; when competence is unavailable, abstain and keep confidence evidence-linked.
@@ Appendix A Lean 4 Proof Listings
The complete Lean 4 formalization is available in the companion artifact (Zenodo DOI listed on the title page). The mechanization consists of 28863 lines across 113 files, with 1252 theorem/lemma statements.
@@ A.1 Lean Handle ID Map
Lean handle entry
AC1DecisionQuotient.ClaimClosure.AtomicCircuitExports.AC1
AC3DecisionQuotient.ClaimClosure.AtomicCircuitExports.AC3
AC4DecisionQuotient.ClaimClosure.AtomicCircuitExports.AC4
AC5DecisionQuotient.ClaimClosure.AtomicCircuitExports.AC5
AC6DecisionQuotient.ClaimClosure.AtomicCircuitExports.AC6
AC8DecisionQuotient.ClaimClosure.AtomicCircuitExports.AC8
AC9DecisionQuotient.ClaimClosure.AtomicCircuitExports.AC9
AC11DecisionQuotient.ClaimClosure.AtomicCircuitExports.AC11
AN1DecisionQuotient.Physics.AssumptionNecessity.no_assumption_free_proof_of_refutable_claim
AN2DecisionQuotient.Physics.AssumptionNecessity.countermodel_violates_some_assumption
AN3DecisionQuotient.Physics.AssumptionNecessity.physical_claim_requires_physical_assumption
AN4DecisionQuotient.Physics.AssumptionNecessity.physical_claim_requires_empirically_justified_physical_assumption
AN5DecisionQuotient.Physics.AssumptionNecessity.strong_physical_no_go_meta
AQ1DecisionQuotient.ClaimClosure.AQ1
AQ2DecisionQuotient.ClaimClosure.AQ2
AQ3DecisionQuotient.ClaimClosure.AQ3
AQ4DecisionQuotient.ClaimClosure.AQ4
AQ5DecisionQuotient.ClaimClosure.AQ5
AQ6DecisionQuotient.ClaimClosure.AQ6
AQ7DecisionQuotient.ClaimClosure.AQ7
AQ8DecisionQuotient.ClaimClosure.AQ8
ARG1PhysicalComplexity.AccessRegime.PhysicalDevice
ARG2PhysicalComplexity.AccessRegime.AccessRegime
ARG3PhysicalComplexity.AccessRegime.RegimeEval
ARG4PhysicalComplexity.AccessRegime.RegimeSample
ARG5PhysicalComplexity.AccessRegime.RegimeProof
ARG6PhysicalComplexity.AccessRegime.RegimeWithCertificate
ARG7PhysicalComplexity.AccessRegime.RegimeEvalOn
ARG8PhysicalComplexity.AccessRegime.RegimeSampleOn
ARG9PhysicalComplexity.AccessRegime.RegimeProofOn
ARG10PhysicalComplexity.AccessRegime.RegimeWithCertificateOn
ARG11PhysicalComplexity.AccessRegime.HardUnderEval
ARG12PhysicalComplexity.AccessRegime.AuditableWithCertificate
ARG13PhysicalComplexity.AccessRegime.certificate_upgrades_regime
ARG14PhysicalComplexity.AccessRegime.certificate_upgrades_regime_on
ARG15PhysicalComplexity.AccessRegime.physical_succinct_certification_hard
ARG16PhysicalComplexity.AccessRegime.certificate_amortizes_hardness
ARG17PhysicalComplexity.AccessRegime.regime_upgrade_with_certificate
ARG18PhysicalComplexity.AccessRegime.regime_upgrade_with_certificate_on
ARG19PhysicalComplexity.AccessRegime.AccessChannelLaw
ARG20PhysicalComplexity.AccessRegime.FiveWayMeet
BA1DecisionQuotient.Physics.BoundedAcquisition.BoundedRegion
BA2DecisionQuotient.Physics.BoundedAcquisition.acquisition_rate_bound
BA3DecisionQuotient.Physics.BoundedAcquisition.acquisitions_are_transitions
BA4DecisionQuotient.Physics.BoundedAcquisition.one_bit_per_transition
BA5DecisionQuotient.Physics.BoundedAcquisition.resolution_reads_sufficient
BA6DecisionQuotient.Physics.BoundedAcquisition.srank_le_resolution_bits
BA7DecisionQuotient.Physics.BoundedAcquisition.energy_ge_srank_cost
BA8DecisionQuotient.Physics.BoundedAcquisition.srank_one_energy_minimum
BA9DecisionQuotient.Physics.BoundedAcquisition.physical_grounding_bundle
BA10DecisionQuotient.Physics.BoundedAcquisition.counting_gap_theorem
BB1DecisionQuotient.BayesianDQ
BB2DecisionQuotient.BayesianDQ.certaintyGain
BB3DecisionQuotient.dq_is_bayesian_certainty_fraction
BB4DecisionQuotient.bayesian_dq_matches_physics_dq
BB5DecisionQuotient.dq_derived_from_bayes
BC1DecisionQuotient.Foundations.counting_nonneg
BC2DecisionQuotient.Foundations.counting_total
BC3DecisionQuotient.Foundations.counting_additive
BC4DecisionQuotient.Foundations.bayes_from_conditional
BC5DecisionQuotient.Foundations.entropy_contraction
BF1DecisionQuotient.certainty_of_not_nondegenerateBelief
BF2DecisionQuotient.nondegenerateBelief_of_uncertaintyForced
BF3DecisionQuotient.forced_action_under_uncertainty
BF4DecisionQuotient.bayes_update_exists_of_nondegenerateBelief
@@
CC3DecisionQuotient.ClaimClosure.system_transfer_licensed_iff_snapshot
@@
CC6DecisionQuotient.ClaimClosure.anchor_query_relation_false_iff
CC7DecisionQuotient.ClaimClosure.anchor_query_relation_true_iff
CC8DecisionQuotient.ClaimClosure.boundaryCharacterized_iff_exists_sufficient_subset
CC9DecisionQuotient.ClaimClosure.bounded_actions_detectable
CC10DecisionQuotient.ClaimClosure.bridge_boundary_represented_family
CC11DecisionQuotient.ClaimClosure.bridge_failure_witness_non_one_step
CC12DecisionQuotient.ClaimClosure.bridge_transfer_iff_one_step_class
CC13DecisionQuotient.ClaimClosure.certified_total_bits_split_core
CC14DecisionQuotient.ClaimClosure.cost_asymmetry_eth_conditional
CC15DecisionQuotient.ClaimClosure.declaredBudgetSlice
CC16DecisionQuotient.ClaimClosure.declaredRegimeFamily_complete
@@
CC17DecisionQuotient.ClaimClosure.declared_physics_no_universal_exact_certifier_core
CC18DecisionQuotient.ClaimClosure.dichotomy_conditional
CC19DecisionQuotient.ClaimClosure.epsilon_admissible_iff_raw_lt_certified_total_core
CC20DecisionQuotient.ClaimClosure.exact_admissible_iff_raw_lt_certified_total_core
@@
CC21DecisionQuotient.ClaimClosure.exact_certainty_inflation_under_hardness_core
CC22DecisionQuotient.ClaimClosure.exact_raw_eq_certified_iff_certainty_inflation_core
CC23DecisionQuotient.ClaimClosure.exact_raw_only_of_no_exact_admissible_core
CC24DecisionQuotient.ClaimClosure.explicit_assumptions_required_of_not_excused_core
CC25DecisionQuotient.ClaimClosure.explicit_state_upper_core
CC26DecisionQuotient.ClaimClosure.hard_family_all_coords_core
CC27DecisionQuotient.ClaimClosure.horizonTwoWitness_immediate_empty_sufficient
CC28DecisionQuotient.ClaimClosure.horizon_gt_one_bridge_can_fail_on_sufficiency
CC29DecisionQuotient.ClaimClosure.information_barrier_opt_oracle_core
CC30DecisionQuotient.ClaimClosure.information_barrier_state_batch_core
CC31DecisionQuotient.ClaimClosure.information_barrier_value_entry_core
CC32DecisionQuotient.ClaimClosure.integrity_resource_bound_for_sufficiency
CC33DecisionQuotient.ClaimClosure.integrity_universal_applicability_core
@@
CC34DecisionQuotient.ClaimClosure.meta_coordinate_irrelevant_of_invariance_on_declared_slice
CC35DecisionQuotient.ClaimClosure.meta_coordinate_not_relevant_on_declared_slice
CC36DecisionQuotient.ClaimClosure.minsuff_collapse_core
CC37DecisionQuotient.ClaimClosure.minsuff_collapse_to_conp_conditional
CC38DecisionQuotient.ClaimClosure.minsuff_conp_complete_conditional
CC39DecisionQuotient.ClaimClosure.no_auto_minimize_of_p_neq_conp
@@
CC40DecisionQuotient.ClaimClosure.no_exact_claim_admissible_under_hardness_core
@@
CC41DecisionQuotient.ClaimClosure.no_exact_claim_under_declared_assumptions_unless_excused_core
CC42DecisionQuotient.ClaimClosure.no_exact_identifier_implies_not_boundary_characterized
CC43DecisionQuotient.ClaimClosure.no_uncertified_exact_claim_core
CC44DecisionQuotient.ClaimClosure.one_step_bridge
CC45DecisionQuotient.ClaimClosure.oracle_lattice_transfer_as_regime_simulation
CC46DecisionQuotient.ClaimClosure.physical_crossover_above_cap_core
CC47DecisionQuotient.ClaimClosure.physical_crossover_core
CC48DecisionQuotient.ClaimClosure.physical_crossover_hardness_core
CC49DecisionQuotient.ClaimClosure.physical_crossover_policy_core
CC50DecisionQuotient.ClaimClosure.process_bridge_failure_witness
CC51DecisionQuotient.ClaimClosure.poseAnchorQuery
CC52DecisionQuotient.ClaimClosure.pose_returns_anchor_query_object
CC53DecisionQuotient.ClaimClosure.posed_anchor_checked_true_implies_truth